# Show-o: One Single Transformer to Unify Multimodal Understanding and Generation

### v4

@@ 5.6 Failure Cases
We provide failure cases of Show-o in multimodal understanding and generation in Fig. 12. The current version of Show-o exhibits challenges in text recognition/generation and object counting. For instance, Show-o struggles to identify the phrase ”closing down” in the left of Fig. 12(a) and is unable to generate the term “mardefly” (as shown left of Fig. 12(b)). This limitation is mainly attributed to the insufficiency of specific data tailored to these scenarios, as our model relies on a limited set of image-text pairs sourced from publicly available datasets and utilizes automatically generated captions. Enriching such kind of data holds promise for addressing these failure modes in Show-o, an aspect that will be explored in the future.
@@ 6 Conclusion
This paper proposed a unified transformer, i.e., Show-o, to unify multimodal understanding and generation. Show-o for the first time unified auto-regressive and (discrete) diffusion modeling that can handle different modalities in distinct ways. Extensive experimental results demonstrated that Show-o is comparable to even better than individual expert models across a wide range of vision-language tasks. This highlighted its potential as a next-generation foundation model.

### v5

@@ 5.6 Failure Cases
We provide failure cases of Show-o in multimodal understanding and generation in Fig. 12. The current version of Show-o sometimes cannot accurately recognize the text and count the object instances and exhibits challenges in generating correct belongings such as skis for each instance. One can observe that Show-o fails to identify the phrase ”closing down” in the left of Fig. 12(a) and is unable to generate the term “mardefly” (as shown left of Fig. 12(b)). This limitation is mainly attributed to the insufficiency of specific data tailored to these scenarios, as our model relies on a limited set of image-text pairs sourced from publicly available datasets and utilizes automatically generated captions. Enriching such kind of data holds promise for addressing these failure modes in Show-o, an aspect that will be explored in the future.
@@ 6 Conclusion
This paper proposed a unified transformer, i.e., Show-o, to unify multimodal understanding and generation. Show-o for the first time unified autoregressive and (discrete) diffusion modeling that can handle different modalities in distinct ways. Extensive experimental results demonstrated that Show-o is comparable to even better than individual expert models across a wide range of vision-language tasks. This highlighted its potential as a next-generation foundation model.